Earth orbits the sun once every 365.25 days.
Find the average angular speed of Earth
about the sun.
Answer in units of rad/s.
Don't round answer
1
Expert's answer
2019-12-17T09:54:43-0500
"\omega=\frac{2\pi}{T}"

"\omega=\frac{2\pi}{(365.25)(24)(3600)}=1.99\cdot 10^{-7}\frac{rad}{s}"